OpenStudy (anonymous):

Luanne bought a sandwich from the vending machine for $1.65. The vending machine takes nickels, dimes, and quarters. She dropped 19 coins into the vending machine to pay for her sandwich. 1. How many of each kind of coin did Luanne use to pay for the sandwich? 2. What strategy did you use to investigate the problem? 3. Are you sure you are correct? How do you know? 4. Is there only one answer to the problem? How do you know?

OpenStudy (mertsj):

There is a piece of information missing.

OpenStudy (anonymous):

this was the only information given...i was able to come up with 2 equations but could get no farther \[n+d+q=19\] and \[.05n+.10d+.25q=1.65\]

OpenStudy (anonymous):

that isn't 165 though

OpenStudy (anonymous):

and thats only 12 coins and i have to have 19 coins total...

OpenStudy (anonymous):

ok i think i have it 2 quarters, 6 dimes, 11 nickels

OpenStudy (anonymous):

thats one solution, i also come up with 14 nickels, 2 dimes, and 3 quarters. so now i know there are multiple solutions but how do i explain the strategy used to solve this problem?

OpenStudy (anonymous):

still not 19 coins though

OpenStudy (anonymous):

do u guys have every thing under controle ?

OpenStudy (anonymous):

ok now i am on a real computer. lets try this \[Q+D+N=19\] \[25Q+10D+5N=165\] now lets get rid of the nickels by multiplying the first equation by -5 and adding to second one

OpenStudy (anonymous):

get \[20Q+5D=70\] or \[D=14-4Q\]

OpenStudy (anonymous):

now we are looking for integer solutions, so you know Q = 1, 2, 3, ... etc

OpenStudy (anonymous):

now get rid of the dimes in a similar fashion

OpenStudy (anonymous):

im with you so far

OpenStudy (anonymous):

btw we really can stop here, because we already know that we have only 3 choices for quarters, 1, 2 and 3 and that determines the number of dimes, so we really don't have to repeat and solve for nickels in terms of quarters

OpenStudy (anonymous):

since we know that \[D=14-4Q\] if there is one quarter, there are ten dimes, and the rest nickels right?

OpenStudy (anonymous):

similarly if there are 2 quarters there are 6 dimes and 11 nickels

OpenStudy (anonymous):

the last one doesnt have 19 coins

OpenStudy (anonymous):

that was wrong. 3 quarters, 2 dimes 14 nickels

OpenStudy (anonymous):

yeah i added wrong sorry.

OpenStudy (anonymous):

ok that sounds better

OpenStudy (anonymous):

so there are three solutions, and they all work. \[\checkmark\]

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

yw